The sum of a student's three scores is 217. If the first is 16 points more than the second, and the sum of the first two is 28 more than twice the third, what was the first score?
Answer by ewatrrr(24785)


Hi
first is 16 points more than the second and sum of the three scores is 217
Let (x+16), x, (217 - (2x +16) represent the 1st,2nd and 3rd scores respectively
Question states***
(x+16) + x = 2[(217 - (2x +16] +28
solving for x
2x + 16 = -4x + 402 + 28
6x = 414
x = 69, the second score, first is 85 (69+16)
CHECKING our Answer***Note 3rd score is 63 (217-69-85)
69 + 85 = 2*63 + 28 = 154